OpenStudy (anonymous):

-2(2z-7y)-z =? Do I do distributive property first? I think I did that and then combined like terms but i keep getting answer incorrect

hartnn (hartnn):

what u got after distributing -2 ??

OpenStudy (anonymous):

-4z+14y-z?

OpenStudy (anonymous):

14y+5z?

hartnn (hartnn):

thats correct now its just -4z-z =... ? 14y remains as it is....

hartnn (hartnn):

-4z-z is not 5z -4z-z = (-4-1)z =... ?

OpenStudy (anonymous):

-4z? Ugh This part confuses me bc do I add or multiple? and if I multiple then it would be 4z

hartnn (hartnn):

its just adding, here its like -ax-bx u get (-a-b)x by factoring out x similarly, -4z-z , by factoring out z, you get (-4-1)z and whats -4-1 =... ?

OpenStudy (anonymous):

14y-5z :)

OpenStudy (unklerhaukus):

\[\color{red}\checkmark\]

hartnn (hartnn):

yes :)